Cita is six years older than bona and bona is twice as old as Albert. Four years from now cita's age will be 8 years less than the sum of albert's and bona's ages. How old is each now
Answer by ankor@dixie-net.com(22740)

Three people; a, b, c
:
Write an equation for each statement
:
Cita is six years older than bona and
c = b + 6
bona is twice as old as Albert.
b = 2a
therefore
c = 2a + 6
Four years from now cita's age will be 8 years less than the sum of albert's and bona's ages.
c + 4 = (a+4) + (b+4) - 8
c + 4 = a + b + 8 - 8
c + 4 == a + b
Replace c with (2a+6); replace b with 2a
(2a+6) + 4 = a + 2a
10 = 3a - 2a
a = 10 yrs is Albert's age
then
b = 2(10)
b = 20 is Bono's age
and
c = 2(10) + 6
c = 26 yr is Cita's age
:
:
Check this in the statement
" Four years from now cita's age will be 8 years less than the sum of albert's and bona's ages. "
26 + 4 = (10+4) + (20+4) - 8
30 = 14 + 24 - 8
30 = 14 + 16